What is the Specific Heat of Copper?

Grade:12th Pass

5 Answers

Raghuvaran varan Chandragiri
11 years ago

386.4 Jkg^-1k^-1

yours katarnak Suresh
11 years ago

Specific heats and molar heat capacities for various substances at 20 C

Substance c in J/gm K c in cal/gm K or
Btu/lb F
Molar C
J/mol K
Aluminum 0.900 0.215 24.3
Bismuth 0.123 0.0294 25.7
Copper 0.386 0.0923 24.5
Brass 0.380 0.092 ...
Gold 0.126 0.0301 25.6
Lead 0.128 0.0305 26.4
Silver 0.233 0.0558 24.9
Tungsten 0.134 0.0321 24.8
Zinc 0.387 0.0925 25.2
Mercury 0.140 0.033 28.3
Alcohol(ethyl) 2.4 0.58 111
Water 4.186 1.00 75.2
Ice (-10 C) 2.05 0.49 36.9
Granite .790 0.19 ...
Glass .84 0.20 ...
Specific Heat Why are the metallic molar specific heats so nearly the same?
Index

Tables

Reference
Tipler
Ch 19
  HyperPhysics***** Thermodynamics Go Back







 

The specific heat of common metals are indicated in the table below:

MetalSpecific Heat - cp
(kJ/kg K) (kcal/kgoC) (Btu/lbmoF)
Aluminum 0.91 0.22 0.22
Antimony 0.21 0.05 0.05
Barium 0.20 0.048 0.048
Beryllium 1.83 0.436 0.436
Bismuth 0.13 0.03 0.03
Cadmium 0.23 0.055 0.055
Calsium 0.63 0.15 0.15
Carbon Steel 0.49 0.12 0.12
Cast Iron 0.46 0.11 0.11
Cesium 0.24 0.057 0.057
Chromium 0.46 0.11 0.11
Cobalt 0.42 0.1 0.1
Copper 0.39 0.092 0.09
Gallium 0.37 0.088 0.088
Germanium 0.32 0.076 0.076
Gold 0.13 0.031 0.03
Hafnium 0.14 0.033 0.033
Indium 0.24 0.057 0.057
Iridium 0.13 0.031 0.31
Iron 0.45 0.108 0.11
Lanthanum 0.195 0.047 0.047
Lead 0.13 0.031 0.03
Lithium 3.57 0.85 0.85
Lutetium 0.15 0.036 0.036
Magnesium 1.05 0.243 0.25
Manganese 0.48 0.114 0.114
Mercury 0.14 0.033 0.03
Molybdenum 0.25 0.06 0.06
Nickel 0.44 0.104 0.10
Niobium (Columbium) 0.27 0.064 0.064
Osmium 0.13 0.031 0.031
Palladium 0.24 0.057 0.057
Platinum 0.13 0.032 0.03
Plutonium 0.13 0.032 0.032
Potassium 0.75 0.180 0.180
Rhenium 0.14 0.033 0.033
Rhodium 0.24 0.058 0.058
Rubidium 0.36 0.086 0.086
Ruthenium 0.24 0.057 0.057
Scandium 0.57 0.14 0.14
Selenium 0.32 0.077 0.077
Silicon 0.71 0.17 0.17
Silver 0.23 0.057 0.057
Sodium 1.21 0.29 0.29
Strontium 0.30 0.072 0.072
Tantalum 0.14 0.034 0.034
Thallium 0.13 0.03 0.03
Thorium 0.13 0.03 0.03
Tin 0.21 0.054 0.05
Titanium 0.54 0.125 0.13
Tungsten 0.13 0.032 0.032
Uranium 0.12 0.028 0.028
Vanadium 0.39 0.116 0.116
Yttrium 0.30 0.072 0.072
Zinc 0.39 0.093 0.09
Zirconium 0.27 0.06 0.06
Wrought Iron 0.50 0.12 0.12
  • 1 J/(kg K) = 2.389x10-4 kcal/(kg oC) = 2.389x10-4 Btu/(lbm oF)
  • 1 kJ/(kg K) = 0.2389 kcal/(kg oC) = 0.2389 Btu/(lbm oF)
  • 1 Btu/(lbm oF) = 4,186.8 J/ (kg K) = 1 kcal/(kg oC)
  • 1 kcal/(kg oC) = 4,186.8 J/ (kg K) = 1 Btu/(lbm oF)

Example - Heating Carbon Steel

2 kg of carbon steel is heated from 20 oC to 100 oC. Specific heat of carbon steel is 0.49 kJ/kg0C and the heat required can be calculated as

dQ = 2 (kg) 0.49 (kJ/kg0C) (100 (oC) - 20 (oC)) 

     = 78.4 (kJ)








ashok kumar
11 years ago

 

Specific heat of copper is   387 joule/kg.ºc

 

KOPPULA ANIL
11 years ago

387 J/Kg. C or  0.0924 Cal/g.C

Jaymit Surve
10 years ago

386.4 j^-1 kg^-1